OpenStudy (anonymous):

y-c/y+c + y+c/y-c = 34/15

jhonyy9 (jhonyy9):

15(y-c)^2 +15(y+c)^2 =34(y+c)(y-c) 15(y^2 -2cy +c^2) +15(y^2 +2cy +c^2) =34(y^2 -c^2) 15(2y^2 +2c^2) =34(y^2 -c^2) 30y^2 +30c^2 =34y^2 -34c^2 64c^2 =4y^2 16c^2 =y^2 y_1,2 = +/- 4c